# To See or Not to See - Fingerprinting Devices in Adversarial Environments Amid Advanced Machine Learning

### v1

@@ IV-B Network Approaches
Features. The features utilized vary depending on the work. Some works gather unique network information from the received packets [76, 77], while other works analyze changes in network traffic caused by devices [3, 78, 79, 67, 80, 81, 82]. CSI:DeSpy captures WiFi camera video bitrate fluctuations as well as CSI data [83].
@@
Mapping (characteristic to a device) and rule-based classifiers can be used, where a dictionary or database is used to store device mappings and data of devices  [76, 3]. Signal processing and statistics techniques such as correlation, peak-finding, statistical testing, and thresholding can be used to detect devices  [78, 67].
Machine learning algorithms can be used to detect hidden devices, ranging from lightweight feature-based classifiers  [80, 83, 79] to more complex deep learning methods  [81].
@@
The receiver (network traffic analyzer) can be implemented on many devices such as smartphones [3, 83, 78, 80], laptops [77], Raspberry Pis [67], or access points [79].
Many transmitting devices are tested such as wireless cameras [77, 83, 78, 82, 67, 80, 81]. In general, many kinds of IoT devices beyond cameras can be tested [76, 79], as long as the device has a wireless connection. This includes devices such as doorbells, bulbs, and smart speakers.
The range for network-based approaches is typically as large as the network range. Only a few approaches are limited in range due to reliance on motion input. For example, for SnoopDog [3], the user must be $<$ 5 meters from the device. For MotionCompass [82], the user must be within the motion detection range of the camera.
The time of detection varies. Some are within a minute [83, 80, 81], while others can be multiple minutes [78, 67] to an hour [76]. Similarly to RF approaches, a shorter detection time is better to limit the negative consequences on users.
A few works consider localization. SnoopDog [3] can localize within a small region of the room. Works such as MotionCompass can localize a device within a couple of centimeters [82, 78].
@@ IV-C Electromagnetic Approaches
Electromagnetic approaches utilize electromagnetic radiation (EMR) from devices in order to detect. We analyze three works in this section. In EarFisher [68], the source of these emissions are from the memory unit. In DeHiRec [84], the emissions are from the ADC. Another work, by Thotla [85] utilizes emissions from a super-regenerative receiver (SRR).
@@
Performance. The receiving devices vary from a BladeRF SDR [68] to an oscilloscope [85]. The transmitters consist of devices such as offline microphones [84], laptops [68] and doorbells [85].
@@ IV-D Thermal Approaches
Thermal approaches attempt to discover a device by detecting the heat dissipation from a device. We present two works: HeatDeCam [69] and See No Evil [86]. In these two works, the devices in question are hidden cameras.
@@
BIAS [87] Impersonate a BLE device during secure connection establishment. RF Packet Address, Role Switch N/A Eval. Board BLE Device
RFAL [88] Identify rogue transmitters via a GAN. RF I/Q Samples I/Q Perturbations D.L. USRP RTL-SDR
GAN Attack [89] Utilize a GAN to produce spoofed signals to fool a fingerprinter. RF I/Q Samples I/Q Perturbations D.L. N/A N/A N/A
GAN Attack [90] Fool RF authentication by adding perturbations in the sent signal. RF Rx Feedback I/Q Perturbations Markov Decision Process ADALM Pluto SDR ADALM Pluto SDR
CNN Attack [5] Applies traditional adversarial attack techniques towards device identification. RF I/Q Feature Map I/Q Perturbations D.L. N/A Signal Generator Signal Receiver
SemperFi [91] Minimize the effect of adversarial GPS spoofing. GPS GPS Signal Adv. Peak Stats. Drone SDR
Hu-Fu [92] RFID authenticator that is resistant to spoofing and replay attacks. RFID RFID Signal Inductive Coupling N/A RFID USRP
Blinding Attack  [93] Blind and confuse actuators to disrupt IoT operation. Network Packet Division Telemetry N/A Wireless Router Home Devices
@@ V Existing Adversarial Attacks
When developing a fingerprinting system, it is important to consider the impact of adversaries on the system, who have their own processing steps ($\Sigma$, $p$ in Figure 3) and $Adversary$ $Models$ to fool the fingerprinter. For example, an attacker can spoof the device. Spoofing is when the adversary attempts to fool the classifier into thinking they are a specific authenticated device (a form of targeted attack), and/or protect themselves from being detected (i.e., a hidden device). An attacker can also conduct an adversarial attack in which the attacker adds perturbations to their signal to fool the classifier, either targeted or untargeted. In either case, the integrity of the fingerprinting system can be put in jeopardy and the system must be designed with security in mind. In the following subsections, we discuss various attacks and defenses for fingerprinting systems. A selection of various works that cover the existing methodologies of spoofing and adversarial attacks are shown in Table III.
@@ V-A Attacks
Input. In most cases, the adversary leverages raw I/Q samples to attack [5, 89, 94, 95, 96, 97]. For BLE spoofing, which involves the spoofing of BLE device identity when pairing, BLE packets are used [98, 87]. In Karunaratne’s work, the attacker receives feedback from the receiver as input to generate adversarial samples [90].
Method of Attack. There are a few methods of attack. It is common for receiver feedback to be used in tandem with a model  [90]. One specific model being used is a generative adversarial network (GAN) [89, 94, 96, 95], which can be used to generate spoofed samples in either a targeted or untargeted setting. In the case of BLE, a model is not used but rather the direct spoofing of identity [98, 87]. Finally, some works have analyzed traditional methods in adversarial machine learning such as FGSM, BIM, and PGD to launch both targeted and untargeted attacks [5, 97].
@@
Another class of attacks attempts to ”blind” the fingerprinter. These works attempt to disable sensors or actuators  [93] or disable or hinder the IoT device  [99].
Fingerprinting Model Attacked. Attacks generally consider attacking deep learning models [5, 89, 94, 95, 96, 97]. Karunaratne attacks a discriminator and an LSTM system [90].
@@
Most of the attacks are simulated or at least conducted partially in a simulated setting. The attacks that leverage traditional methods in adversarial machine learning generally do not consider over-the-air methods and only consider attacking the received input [5, 97].
Since many of the attacks were simulated, those works do not mention the distance that this attack was successful. For the works that are conducted in real settings, the range of the wireless medium is assumed [98, 87, 90] or within the wireless range  [93, 99].
The time to conduct the attack is generally tied to the time it takes to train the model [5, 89, 94]. In the case of BLE attacks, the attack can be as quick as seconds or within pairing time [98, 87].
In many cases, physical access to the devices is needed in order to place an adversarial receiver close to the receiver [89] or to access the samples for adversarial machine learning techniques [5, 97]. In some works, the adversary needs to be close enough to the victim devices in order to capture their signals [94, 98, 87] or the receiver’s signals [90].
@@ V-B Defenses
Input. Defenses towards spoofing or adversarial attacks receive the same kind of signal (I/Q samples, packets, etc.), no matter if the signal is trusted or adversarial. I/Q samples are used in many works [95, 96, 97, 88]. Features of the packet itself can be used as well, such as CSI and RSS [100, 101, 102]. In other cases, the raw signal itself is used, such as in GPS signals [91] or RFID signals [92].
Method of Defense. The prescribed defense for spoofing and adversarial attacks depends on the signal being sent and the motivation of the defender. For example, a defender can attempt to prevent attacks from happening in the first place. By using techniques such as adversarial training  [96, 95, 97] or perturbation subtracting [95], in addition to designing the authentication scheme more securely  [92], a defender can train the classifier to be robust against adversarial samples.
A defender can also design a defense to detect adversaries. Anomaly detectors can be designed to separate authenticated devices from foreign devices. These can be designed by analyzing deviations  [91] or by designing more complex models  [101, 102, 88].
@@
In most cases, the setup exists in a physical testbed [92, 100, 88, 101, 102]. The distance between the detector and the adversary is generally large. The range can be as small as 10 feet [88] or as large as 4.1 km [91]. Works that leverage RSS have a range in the tens of meters [101, 102].
The time of defense is generally quick and can be a few seconds [91, 92, 100] or a few minutes [101, 102].

### v2

@@ IV-B Network Approaches
Features. The features utilized vary depending on the work. Some works gather unique network information from the received packets [saidi2020haystack, siby2017iotscanner], while other works analyze changes in network traffic caused by devices [singh2021always, heo2022there, ortiz2019devicemien, liu2018detecting, cheng2018dewicam, wu2019you, he2021motioncompass]. CSI:DeSpy captures WiFi camera video bitrate fluctuations as well as CSI data [salman2022csi].
@@
Mapping (characteristic to a device) and rule-based classifiers can be used, where a dictionary or database is used to store device mappings and data of devices  [saidi2020haystack, singh2021always]. Signal processing and statistics techniques such as correlation, peak-finding, statistical testing, and thresholding can be used to detect devices  [heo2022there, liu2018detecting].
Machine learning algorithms can be used to detect hidden devices, ranging from lightweight feature-based classifiers  [cheng2018dewicam, salman2022csi, ortiz2019devicemien] to more complex deep learning methods  [wu2019you].
@@
The receiver (network traffic analyzer) can be implemented on many devices such as smartphones [singh2021always, salman2022csi, heo2022there, cheng2018dewicam], laptops [siby2017iotscanner], Raspberry Pis [liu2018detecting], or access points [ortiz2019devicemien].
Many transmitting devices are tested such as wireless cameras [siby2017iotscanner, salman2022csi, heo2022there, he2021motioncompass, liu2018detecting, cheng2018dewicam, wu2019you]. In general, many kinds of IoT devices beyond cameras can be tested [saidi2020haystack, ortiz2019devicemien], as long as the device has a wireless connection. This includes devices such as doorbells, bulbs, and smart speakers.
The range for network-based approaches is typically as large as the network range. Only a few approaches are limited in range due to reliance on motion input. For example, for SnoopDog [singh2021always], the user must be << 5 meters from the device. For MotionCompass [he2021motioncompass], the user must be within the motion detection range of the camera.
The time of detection varies. Some are within a minute [salman2022csi, cheng2018dewicam, wu2019you], while others can be multiple minutes [heo2022there, liu2018detecting] to an hour [saidi2020haystack]. Similarly to RF approaches, a shorter detection time is better to limit the negative consequences on users.
A few works consider localization. SnoopDog [singh2021always] can localize within a small region of the room. Works such as MotionCompass can localize a device within a couple of centimeters [he2021motioncompass, heo2022there].
@@ IV-C Electromagnetic Approaches
Electromagnetic approaches utilize electromagnetic radiation (EMR) from devices in order to detect. We analyze three works in this section. In EarFisher [shen2021earfisher], the source of these emissions are from the memory unit. In DeHiRec [zhou2022dehirec], the emissions are from the ADC. Another work, by Thotla [thotla2013detection] utilizes emissions from a super-regenerative receiver (SRR).
@@
Performance. The receiving devices vary from a BladeRF SDR [shen2021earfisher] to an oscilloscope [thotla2013detection]. The transmitters consist of devices such as offline microphones [zhou2022dehirec], laptops [shen2021earfisher] and doorbells [thotla2013detection].
@@ IV-D Thermal Approaches
Thermal approaches attempt to discover a device by detecting the heat dissipation from a device. We present two works: HeatDeCam [yu2022heatdecam] and See No Evil [zuniga2022see]. In these two works, the devices in question are hidden cameras.
@@
BIAS [antonioli2020bias] Impersonate a BLE device during secure connection establishment. RF Packet Address, Role Switch N/A Eval. Board BLE Device
RFAL [roy2019rfal] Identify rogue transmitters via a GAN. RF I/Q Samples I/Q Perturbations D.L. USRP RTL-SDR
GAN Attack [shi2020generative] Utilize a GAN to produce spoofed signals to fool a fingerprinter. RF I/Q Samples I/Q Perturbations D.L. N/A N/A N/A
GAN Attack [karunaratne2021penetrating] Fool RF authentication by adding perturbations in the sent signal. RF Rx Feedback I/Q Perturbations Markov Decision Process ADALM Pluto SDR ADALM Pluto SDR
CNN Attack [bao2021threat] Applies traditional adversarial attack techniques towards device identification. RF I/Q Feature Map I/Q Perturbations D.L. N/A Signal Generator Signal Receiver
SemperFi [sathaye2021semperfi] Minimize the effect of adversarial GPS spoofing. GPS GPS Signal Adv. Peak Stats. Drone SDR
Hu-Fu [wang2018towards] RFID authenticator that is resistant to spoofing and replay attacks. RFID RFID Signal Inductive Coupling N/A RFID USRP
Blinding Attack  [oconnor2019blinded] Blind and confuse actuators to disrupt IoT operation. Network Packet Division Telemetry N/A Wireless Router Home Devices
@@ V Existing Adversarial Attacks
When developing a fingerprinting system, it is important to consider the impact of adversaries on the system, who have their own processing steps ($\Sigma$, $p$ in Figure 3) and AdversaryAdversary ModelsModels to fool the fingerprinter. For example, an attacker can spoof the device. Spoofing is when the adversary attempts to fool the classifier into thinking they are a specific authenticated device (a form of targeted attack), and/or protect themselves from being detected (i.e., a hidden device). An attacker can also conduct an adversarial attack in which the attacker adds perturbations to their signal to fool the classifier, either targeted or untargeted. In either case, the integrity of the fingerprinting system can be put in jeopardy and the system must be designed with security in mind. In the following subsections, we discuss various attacks and defenses for fingerprinting systems. A selection of various works that cover the existing methodologies of spoofing and adversarial attacks are shown in Table III.
@@ V-A Attacks
Input. In most cases, the adversary leverages raw I/Q samples to attack [bao2021threat, shi2020generative, xu2022colluding, bahramali2021robust, merchant2019securing, xu2023adversarial]. For BLE spoofing, which involves the spoofing of BLE device identity when pairing, BLE packets are used [ai2022blacktooth, antonioli2020bias]. In Karunaratne’s work, the attacker receives feedback from the receiver as input to generate adversarial samples [karunaratne2021penetrating].
Method of Attack. There are a few methods of attack. It is common for receiver feedback to be used in tandem with a model  [karunaratne2021penetrating]. One specific model being used is a generative adversarial network (GAN) [shi2020generative, xu2022colluding, merchant2019securing, bahramali2021robust], which can be used to generate spoofed samples in either a targeted or untargeted setting. In the case of BLE, a model is not used but rather the direct spoofing of identity [ai2022blacktooth, antonioli2020bias]. Finally, some works have analyzed traditional methods in adversarial machine learning such as FGSM, BIM, and PGD to launch both targeted and untargeted attacks [bao2021threat, xu2023adversarial].
@@
Another class of attacks attempts to ”blind” the fingerprinter. These works attempt to disable sensors or actuators  [oconnor2019blinded] or disable or hinder the IoT device  [fu2022iot].
Fingerprinting Model Attacked. Attacks generally consider attacking deep learning models [bao2021threat, shi2020generative, xu2022colluding, bahramali2021robust, merchant2019securing, xu2023adversarial]. Karunaratne attacks a discriminator and an LSTM system [karunaratne2021penetrating].
@@
Most of the attacks are simulated or at least conducted partially in a simulated setting. The attacks that leverage traditional methods in adversarial machine learning generally do not consider over-the-air methods and only consider attacking the received input [bao2021threat, xu2023adversarial].
Since many of the attacks were simulated, those works do not mention the distance that this attack was successful. For the works that are conducted in real settings, the range of the wireless medium is assumed [ai2022blacktooth, antonioli2020bias, karunaratne2021penetrating] or within the wireless range  [oconnor2019blinded, fu2022iot].
The time to conduct the attack is generally tied to the time it takes to train the model [bao2021threat, shi2020generative, xu2022colluding]. In the case of BLE attacks, the attack can be as quick as seconds or within pairing time [ai2022blacktooth, antonioli2020bias].
In many cases, physical access to the devices is needed in order to place an adversarial receiver close to the receiver [shi2020generative] or to access the samples for adversarial machine learning techniques [bao2021threat, xu2023adversarial]. In some works, the adversary needs to be close enough to the victim devices in order to capture their signals [xu2022colluding, ai2022blacktooth, antonioli2020bias] or the receiver’s signals [karunaratne2021penetrating].
@@ V-B Defenses
Input. Defenses towards spoofing or adversarial attacks receive the same kind of signal (I/Q samples, packets, etc.), no matter if the signal is trusted or adversarial. I/Q samples are used in many works [bahramali2021robust, merchant2019securing, xu2023adversarial, roy2019rfal]. Features of the packet itself can be used as well, such as CSI and RSS [liu2019real, chen2007detecting, sheng2008detecting]. In other cases, the raw signal itself is used, such as in GPS signals [sathaye2021semperfi] or RFID signals [wang2018towards].
Method of Defense. The prescribed defense for spoofing and adversarial attacks depends on the signal being sent and the motivation of the defender. For example, a defender can attempt to prevent attacks from happening in the first place. By using techniques such as adversarial training  [merchant2019securing, bahramali2021robust, xu2023adversarial] or perturbation subtracting [bahramali2021robust], in addition to designing the authentication scheme more securely  [wang2018towards], a defender can train the classifier to be robust against adversarial samples.
A defender can also design a defense to detect adversaries. Anomaly detectors can be designed to separate authenticated devices from foreign devices. These can be designed by analyzing deviations  [sathaye2021semperfi] or by designing more complex models  [chen2007detecting, sheng2008detecting, roy2019rfal].
@@
In most cases, the setup exists in a physical testbed [wang2018towards, liu2019real, roy2019rfal, chen2007detecting, sheng2008detecting]. The distance between the detector and the adversary is generally large. The range can be as small as 10 feet [roy2019rfal] or as large as 4.1 km [sathaye2021semperfi]. Works that leverage RSS have a range in the tens of meters [chen2007detecting, sheng2008detecting].
The time of defense is generally quick and can be a few seconds [sathaye2021semperfi, wang2018towards, liu2019real] or a few minutes [chen2007detecting, sheng2008detecting].